OpenStudy (anonymous):

Evaluate the expression for x = 4 and y = 5. 6(x2 - y) - y2

OpenStudy (anonymous):

INeed HeLp !

OpenStudy (anonymous):

what is required? ..... substitute , multiply . end!

OpenStudy (anonymous):

6(x^2-y)-y^2=6(16-5)-25=6*9-25=54-25=29 if anyone knows a better solution... post it :)